# Sprecher Networks: A Parameter-Efficient Kolmogorov-Arnold Architecture

### v1

@@ 7.4 Theoretical domain computation
When the domain of a $\Phi^{(\ell)}$ spline changes, new knot locations are established (typically uniformly spaced across the new computed domain). The original spline is treated as a continuous function and evaluated at these new knot locations to yield updated knot coefficients (for piecewise-linear $\Phi^{(\ell)}$, query points are clamped to the previous domain so any new knot outside the old domain uses the nearest boundary value; for cubic/PCHIP $\Phi^{(\ell)}$, the previous spline’s linear extrapolation is used outside the old domain). This effectively “resamples” the learned shape onto the new domain: it preserves the values of the old spline at the new knot locations that lie within the previous domain (and therefore preserves the function exactly on the overlap whenever the old spline is linear on each new knot interval, e.g. when the new knot set contains the old knots). For piecewise-linear $\Phi^{(\ell)}$, this procedure is exact whenever the new knot set contains the old knots; for cubic (PCHIP) $\Phi^{(\ell)}$, it matches the old spline at the resampled knots but can (slightly) change derivatives between knots, while still avoiding abrupt resets. Optionally, one may additionally apply a cheap post-update refresh that adjusts knot locations without resampling coefficients; this is not function-preserving (it amounts to a mild horizontal rescaling), so it is best reserved for small domain changes where the primary goal is to keep bound propagation synchronized after parameter updates.
@@
The ability to compute provably safe (typically conservative) domain and range bounds through interval arithmetic provides several practical benefits: (i) it enables theoretically grounded initialization without arbitrary hyperparameters, (ii) it can help diagnose training issues by detecting when values fall outside expected ranges, (iii) it promotes numerical stability by encouraging spline evaluations to remain within their well-resolved knot regions (with explicit extension/extrapolation behavior outside the knot domain), and (iv) it allows for adaptive domain adjustments that account for lateral mixing dynamics. These benefits distinguish Sprecher Networks from architectures with less structured internal representations.
@@ 8 Empirical demonstrations and case studies
While comprehensive benchmarking remains future work, we provide initial empirical demonstrations to illustrate the feasibility and characteristics of SNs, including the impact of lateral mixing connections.
@@
We evaluate Sprecher Networks on synthetic supervised regression tasks and on a physics-informed Poisson problem.
@@
$\Delta f(\mathbf{x})=g(\mathbf{x})\quad\text{for }\mathbf{x}\in\Omega,\qquad f(\mathbf{x})=0\quad\text{for }\mathbf{x}\in\partial\Omega,$
@@
so that $g=\Delta u$ is known analytically. We train with collocation sets $S_{\mathrm{int}}\subset\Omega$ and $S_{\partial}\subset\partial\Omega$ using
$\mathcal{L}\;=\;\frac{1}{|S_{\mathrm{int}}|}\sum_{\mathbf{x}\in S_{\mathrm{int}}}\frac{\bigl(\Delta f(\mathbf{x})-g(\mathbf{x})\bigr)^{2}}{\mathbb{E}_{\mathbf{x}\in S_{\mathrm{int}}}[g(\mathbf{x})^{2}]}\;+\;\frac{1}{|S_{\partial}|}\sum_{\mathbf{x}\in S_{\partial}}f(\mathbf{x})^{2}.$
@@
Model variants for the PINN benchmark. To isolate the spline parameterization itself, we compare a barebones Sprecher Network (no lateral mixing, no residuals, no normalization, and fixed spline domains; i.e. no domain updates) against a barebones KAN baseline with spline-only edge activations (no SiLU/base activation and no grid updates). Because the Poisson residual involves second derivatives, both models use shape-preserving cubic (PCHIP) splines in this benchmark (piecewise-linear splines yield $\Delta f_{\theta}=0$ almost everywhere). For the Sprecher Network, inputs are linearly rescaled from $1,1]^{2}$ to $0,1]^{2}$ before the first spline evaluation, matching the $[0,1]$ spline-domain convention used for SNs in this work; the KAN baseline operates directly on $1,1]^{2}$. In both cases, derivatives are taken with respect to the original coordinates via automatic differentiation (for SN this is through the rescaling map).
Optimization and evaluation. We train in torch.float64 using Adam with $10^{-3}$ learning rate, cosine learning rate decay to $10^{-5}$, and gradient-norm clipping 1.0. We report two parameter budgets: an $\approx$1200-parameter model trained for 5,000 epochs (mean $\pm$ std over 5 seeds), and an $\approx$3000-parameter model trained for 20,000 epochs (1 seed). We evaluate both models on a uniform $101\times 101$ grid over $\Omega=[-1,1]^{2}$ and report: (i) PDE residual MSE (interior grid points), (ii) boundary MSE (boundary grid points), and (iii) $L^{2}$ MSE against the manufactured solution (full grid).
@@ 8.1 Basic function approximation
Figure 8: Example of SN approximating the scalar 2D target function $z=f(x,y)=\exp(\sin(11x))+3y+4\sin(8y)$ on $(x,y)\in[0,1]^{2}$ using architecture $2\to[5,8,5]\to 1$ (3 blocks). Top row: Learned spline functions for each block — monotonic splines $\phi^{(\ell)}$ (cyan) and general splines $\Phi^{(\ell)}$ (magenta). Bottom row: Comparison between the target function surface (left) and the network approximation (right). This run uses cyclic lateral mixing (default); the learned outer splines $\Phi^{(\ell)}$ can be highly oscillatory/jagged even when the overall fit is accurate, so the per-layer spline plots should be interpreted qualitatively.
@@ 8.2 Impact of lateral mixing
For many scalar-output regression tasks with moderate widths, lateral mixing yields only modest gains. However, scalar-output SNs can still benefit substantially in regimes where the shared-weight constraint induces strong symmetry across channels (e.g., very wide shallow networks), where lateral mixing provides lightweight cross-channel interaction that can break optimization plateaus (see the $2\to[120]\to 1$ example above). Note that the final summation is a fixed linear readout after applying $\Phi$ channelwise; it does not provide pre-$\Phi$ cross-channel communication, so lateral mixing remains a distinct (and sometimes crucial) mechanism even when the network output is scalar.
@@
Unless otherwise stated (e.g., Figure 10), each head-to-head uses exactly $4000$ training epochs for both architectures. Both models use the same linear residual style: a learned scalar gate when $d_{\mathrm{in}}=d_{\mathrm{out}}$, and a learned projection matrix $W\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ when $d_{\mathrm{in}}\neq d_{\mathrm{out}}$. KANs use shape-preserving cubic Hermite (PCHIP; $\deg=3$) edge splines. SNs use domain warmup early in training (300–400 epochs depending on the task; relative domain safety margin $0.01$) and then freeze the learned domains; lateral mixing is disabled (i.e., $\tau^{(\ell)}=0$) in these parity runs. We match parameter counts by choosing the KAN spline size $K$ (the number of learnable knot values / spline coefficients per univariate edge spline) to be as close as possible (preferring $\leq$) to the SN parameter count, and we keep batch-normalization (BN) semantics aligned at test time with the training semantics. Seeds, datasets, and train/test splits are identical across models. Training is full-batch (no mini-batching): each epoch uses the entire training set. Primary metrics are test set RMSE or MSE (reported per task below); for multi-head monotone problems we also report a monotonicity-violation rate and a correlation-structure error.

### v2

@@ 7.5 Theoretical domain computation
When the domain of a $\Phi^{(\ell)}$ spline changes, new knot locations are established (typically uniformly spaced across the new computed domain). The original spline is treated as a continuous function and evaluated at these new knot locations to yield updated knot coefficients (for piecewise-linear $\Phi^{(\ell)}$, query points are clamped to the previous domain so any new knot outside the old domain uses the nearest boundary value; for cubic/PCHIP $\Phi^{(\ell)}$, out-of-domain query points are evaluated via the spline’s linear extrapolation determined by the endpoint derivative). This effectively “resamples” the learned shape onto the new domain: it preserves the values of the old spline at the new knot locations that lie within the previous domain (and therefore preserves the function exactly on the overlap whenever the old spline is linear on each new knot interval, e.g. when the new knot set contains the old knots). For piecewise-linear $\Phi^{(\ell)}$, this procedure is exact whenever the new knot set contains the old knots; for cubic (PCHIP) $\Phi^{(\ell)}$, it matches the old spline at the resampled knots but can (slightly) change derivatives between knots, while still avoiding abrupt resets. Optionally, one may additionally apply a cheap post-update refresh that adjusts knot locations without resampling coefficients; this is not value-preserving (it amounts to a mild horizontal rescaling), so it is best reserved for small domain changes where the primary goal is to keep bound propagation synchronized after parameter updates.
@@
The ability to compute sound (typically conservative) domain and range bounds through interval arithmetic (provably sound in the absence of normalization; with BatchNorm in training mode we use the heuristic bounds from Section 3.3) provides several practical benefits: (i) it enables theoretically grounded initialization without arbitrary hyperparameters, (ii) it can help diagnose training issues by detecting when values fall outside expected ranges, (iii) it promotes numerical stability by encouraging spline evaluations to remain within their well-resolved knot regions (with explicit extension/extrapolation behavior outside the knot domain), and (iv) it allows for adaptive domain adjustments that account for lateral mixing dynamics. These benefits distinguish Sprecher Networks from architectures with less structured internal representations.
@@ 8 Empirical demonstrations and case studies
We report a set of empirical studies designed to probe both performance and scaling behavior of SNs, with controlled comparisons to MLP and KAN baselines.
@@
We evaluate Sprecher Networks on synthetic supervised regression tasks, on a physics-informed Poisson problem, and on the Fashion-MNIST image classification benchmark.
Fashion-MNIST classification benchmark.
To assess SNs on a standard real-world benchmark, we train on Fashion-MNIST [17], a 10-class image classification dataset of $28\times 28$ grayscale images with 60,000 training and 10,000 test examples (784 input dimensions after flattening; we use the standard split). We use cross-entropy loss and report test accuracy (mean $\pm$ std over 10 seeds). Inputs are scaled to $[0,1]$ (raw pixel intensities rescaled by $1/255$) before being fed to both SN and MLP models. Unless stated otherwise, we train with AdamW (learning rate $10^{-3}$, batch size 64, weight decay $10^{-6}$) and clip gradients to max-norm 1. For the MLP baselines we use ReLU activations and no BatchNorm layers. For SNs we use cubic (PCHIP) splines with 60 knots for the inner maps $\phi$; for the outer maps $\Phi$ we use 60 knots for the shallow models and 180 knots for the 25-layer model (Table 6). We apply BatchNorm after each block (except the first) and use residual connections in all SN Fashion-MNIST runs (linear-projection residuals for the $[12,11,12]$ models; cyclic residuals for the constant-width 25-layer model). We enable cyclic lateral mixing in all SN Fashion-MNIST runs, and we update (theoretical) spline domains before each minibatch during both training and evaluation (resampling the knot values of $\Phi$ accordingly). When reporting test accuracy, we evaluate BatchNorm using per-batch statistics while freezing its running-statistic buffers. Table 6 compares SNs against an MLP baseline under matched parameter budgets.
Table 6: Fashion-MNIST classification results. Test accuracy (mean $\pm$ std over 10 seeds). The shallow SN uses linear residual projections; the 25-layer SN uses cyclic residuals and demonstrates stable training at 25 hidden layers when equipped with residual connections and BatchNorm.
Model Architecture Params Epochs Best Test Accuracy (%)
SN $784\to[12,11,12]\to 10$ 11,244 20 $85.09\pm 0.18$
SN $784\to[12,11,12]\to 10$ 11,244 200 $86.10\pm 0.28$
SN (25 layers) $784\to[40]^{25}\to 10$ 12,008 20 $85.90\pm 0.24$
MLP $784\to[15,20]\to 10$ 12,305 20 $86.27\pm 0.09$
MLP $784\to[15,20]\to 10$ 12,305 200 $86.76\pm 0.15$
At this parameter budget, the shallow SN is competitive with the MLP baseline. More importantly for trainability, the constant-width 25-layer SN optimizes stably across seeds once equipped with cyclic residual connections and normalization, reaching $85.90\pm 0.24\%$ test accuracy. This suggests that deep stacks of Sprecher blocks can be trained reliably, although closing the remaining gap to dense MLPs (and understanding when depth helps) remains an open direction.
@@
$\Delta u(\mathbf{x})=g(\mathbf{x})\quad\text{for }\mathbf{x}\in\Omega,\qquad u(\mathbf{x})=0\quad\text{for }\mathbf{x}\in\partial\Omega,$
@@
so that $g=\Delta u$ is known analytically. We train a network $u_{\theta}$ with collocation sets $S_{\mathrm{int}}\subset\Omega$ and $S_{\partial}\subset\partial\Omega$ by minimizing
$\mathcal{L}\;=\;\frac{1}{|S_{\mathrm{int}}|}\sum_{\mathbf{x}\in S_{\mathrm{int}}}\frac{\bigl(\Delta u_{\theta}(\mathbf{x})-g(\mathbf{x})\bigr)^{2}}{\mathbb{E}_{\mathbf{x}\in S_{\mathrm{int}}}[g(\mathbf{x})^{2}]}\;+\;\frac{1}{|S_{\partial}|}\sum_{\mathbf{x}\in S_{\partial}}u_{\theta}(\mathbf{x})^{2}.$
@@
Model variants for the PINN benchmark. To isolate the spline parameterization itself, we compare a barebones Sprecher Network (no lateral mixing, no residuals, no normalization, and fixed spline domains; i.e. no domain updates) against a barebones KAN baseline with spline-only edge activations (no SiLU/base activation and no grid updates). Because the Poisson residual involves second derivatives, both models use shape-preserving cubic (PCHIP) splines in this benchmark (piecewise-linear splines yield $\Delta u_{\theta}=0$ almost everywhere). For the Sprecher Network, inputs are linearly rescaled from $1,1]^{2}$ to $0,1]^{2}$ before the first spline evaluation via the affine map $\rho:\Omega\to[0,1]^{2}$ defined component-wise by $\rho(\mathbf{x})=(\mathbf{x}+\mathbf{1})/2$, matching the $[0,1]$ spline-domain convention used for SNs in this work; the KAN baseline operates directly on $1,1]^{2}$. In both cases, derivatives are taken with respect to the original coordinates $\mathbf{x}$ via automatic differentiation (for SN this includes the Jacobian of $\rho$ by the chain rule).
Optimization and evaluation. We train in double precision (float64) using Adam with $10^{-3}$ learning rate, cosine learning rate decay to $10^{-5}$, and gradient-norm clipping 1.0. We report two parameter budgets: an $\approx$1200-parameter model trained for 5,000 epochs (mean $\pm$ std over 5 seeds), and an $\approx$3000-parameter model trained for 20,000 epochs (1 seed). We evaluate both models on a uniform $101\times 101$ grid over $\Omega=[-1,1]^{2}$ and report: (i) PDE residual MSE (interior grid points), (ii) boundary MSE (boundary grid points), and (iii) $L^{2}$ MSE against the manufactured solution (full grid).
@@ 8.1 Basic function approximation
Figure 10: Example of SN approximating the scalar 2D target function $z=f(x,y)=\exp(\sin(11x))+3y+4\sin(8y)$ on $(x,y)\in[0,1]^{2}$ using architecture $2\to[5,8,5]\to 1$ (3 blocks). Top row: Learned spline functions for each block — monotonic splines $\phi^{(\ell)}$ (cyan) and general splines $\Phi^{(\ell)}$ (magenta). Bottom row: Comparison between the target function surface (left) and the network approximation (right). This run uses cyclic lateral mixing; the learned outer splines $\Phi^{(\ell)}$ can be highly oscillatory/jagged even when the overall fit is accurate, so the per-layer spline plots should be interpreted qualitatively.
@@ 8.2 Impact of lateral mixing
For many scalar-output regression tasks with moderate widths, lateral mixing yields only modest gains. However, as discussed in Remark 3.6, scalar-output SNs can benefit substantially in regimes where the shared-weight constraint induces strong symmetry across channels (e.g., very wide shallow networks), where lateral mixing provides lightweight cross-channel interaction that breaks optimization plateaus. Note that the final summation is a fixed linear readout after applying $\Phi$ channelwise; it does not provide pre-$\Phi$ cross-channel communication, so lateral mixing remains a distinct (and sometimes crucial) mechanism even when the network output is scalar.
@@
Unless otherwise stated (e.g., Figure 12), each head-to-head uses exactly $4000$ training epochs for both architectures. All benchmark results in this subsection use the following comparison setting:
  • Barebones comparisons8.3.1): Both SN and KAN use no residual connections, no normalization layers, and (for SN) no lateral mixing. SNs use piecewise-linear (PWL) splines with dynamic spline-domain updates during the first 10% of training (400/4000 epochs), then frozen. KANs use cubic Hermite (PCHIP) splines with fixed uniform knot grids—no grid updates or adaptive knot relocation. This isolates the core architectural differences.
We match parameter counts by choosing the KAN spline basis size $K$ (the number of learnable knot values per edge spline) to be as close as possible to the SN parameter count (allowing small over/under-shoot when exact parity is not attainable). Seeds, datasets, and train/test splits are identical across models. Training uses either full-batch or mini-batch updates depending on the benchmark; when mini-batching is used, SN and KAN are trained on identical mini-batches in the same order. Primary metric is test-set RMSE (reported per task below; for vector-valued outputs, RMSE is computed over all test samples and output coordinates). For Benchmark 3 we report the best (minimum) test RMSE attained during training (a non-standard choice since it uses the test set for checkpoint selection); for Benchmark 8 (which includes a validation split) we report the test RMSE at the checkpoint with best validation RMSE.